OpenStudy (xxlyh):

Fan and Medal! An average clementine is shaped like a sphere and has a diameter of 3.5 inches. What is the volume of the fruit? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (xxlyh):

I chose C^^

OpenStudy (mrnood):

are you confident with that? how did you calculate it?

OpenStudy (xxlyh):

I calculated it again and I'm kinda thinking it's A sorry TT

OpenStudy (xxlyh):

d=3.5 r=d/2 r=3.5/2

OpenStudy (mrnood):

just take time and care it is a simple equation : V = 4 pi r^3 /3

OpenStudy (mrnood):

btw - I only asked if you were confident (in case you just guessed)

OpenStudy (mrnood):

so - you are correct r= 1.75 (i.e. 3.5/2) check that in 4 * 3.14 * r^3/3

OpenStudy (xxlyh):

I got 21..34 ^

OpenStudy (mrnood):

I suggest you try again

OpenStudy (mrnood):

4 * 3.14 * r^3/3 r = 1.75

OpenStudy (xxlyh):

21.98 is what I'm getting

OpenStudy (mrnood):

21.98 is not one of th eanswer options - so it is wrong if you are using a calculator try doing th e 'cubed 'f irst: r^3 * 4 * 3.14 /3 r = 1.75

OpenStudy (xxlyh):

22.4379166667

OpenStudy (xxlyh):

is what I got^

OpenStudy (xxlyh):

so am I correct when I say C is the answer? haha

OpenStudy (xxlyh):

@MrNood